# ReCoRe: Regularized Contrastive Representation Learning of World Model

### v1

@@ 4.5 Ablation Study
The results for contribution of constrative learning and regularization in the form of intervention invariant auxiliary tasks are also shown in Table 1. The standard formulation of contrastive learning does not use reconstruction loss [3, 8, 20, 16]. Since model-based RL does not optimize the representation learning and controller jointly, contrastive loss collapses. Hence, to validate our proposal of intervention invariant depth reconstruction as a regularizer, we have done experiments without depth reconstruction (ReCoRe - D). We can see in Table 1 that in a reasonably complex pixel-based control task, ReCoRe is not able to learn the meaningful control without reconstruction task. Further, reconstructing RGB image (I) instead of depth (D), i.e. ReCoRe - D + I, slightly improves the results, but is still approximately three time worse than the proposed ReCoRe.
@@ 5 Conclusion
We proposed a method to learn a World Model with invariant features, ReCoRe. These invariant features are learnt by minimizing contrastive loss between content invariance interventions of the observation. Hence, we proposed auxiliary task as a regularizer, which is invariant to the proposed data augmentation techniques. ReCoRe significantly outperforms the state-of-the-art models on OoD generalization, sim-to-real transfer and sample efficiency measures. As such, ReCoRe is a new state of the art in model-based RL for sample efficiency in OoD generalization. Finally, we note that our framework can be applied to other tasks and the design of interventions and invariant auxiliary losses will become an interesting research problem.
Appendix
Appendix A Qualitative Results of ReCoRe
@@ Appendix B Hyper Parameters
Appendix C iGibson 1.0 Training and Evaluation Splits
@@
concrete Train [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Test [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
fabric Train [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Test [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
planks Train [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Test [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
wood Train [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Test [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]

### v2

@@ 4.5 Ablation Study
The contribution of contrastive learning and regularization in the form of intervention invariant auxiliary tasks are also shown in Table 1. The standard formulation of contrastive learning does not use reconstruction loss [3, 8, 20, 16]. Since model-based RL does not optimize the representation learning and controller jointly, contrastive loss collapses. Hence, to validate our proposal of intervention invariant depth reconstruction as a regularizer, we have done experiments without depth reconstruction (ReCoRe - D). We can see in Table 1 that in a reasonably complex pixel-based control task, ReCoRe is not able to learn meaningful control without the reconstruction task. Further, reconstructing RGB image (I) instead of depth (D), i.e. ReCoRe - D + I, slightly improves the results over ReCoRe - D, but is still approximately three times worse than the proposed ReCoRe.
@@ 5 Conclusion
We proposed a method to learn a World Model with invariant features, ReCoRe. These invariant features are learned by minimizing contrastive loss between content invariance interventions of the observation. Hence, we proposed an auxiliary task as a regularizer, which is invariant to the proposed data augmentation techniques. ReCoRe significantly outperforms the state-of-the-art models on OoD generalization, sim-to-real transfer and sample efficiency measures. As such, ReCoRe is a new state of the art in model-based RL for sample efficiency in OoD generalization. Finally, we note that our framework can be applied to other tasks and the design of interventions and invariant auxiliary losses will become an interesting research problem.
Supplementary Material
1 Qualitative Results of ReCoRe
@@ 2 Hyper Parameters
3 iGibson 1.0 Training and Evaluation Splits
@@
concrete Train Refer to caption Refer to caption Refer to caption
Test Refer to caption Refer to caption Refer to caption
fabric Train Refer to caption Refer to caption Refer to caption
Test Refer to caption Refer to caption Refer to caption
planks Train Refer to caption Refer to caption Refer to caption
Test Refer to caption Refer to caption Refer to caption
wood Train Refer to caption Refer to caption Refer to caption
Test Refer to caption Refer to caption Refer to caption